OpenStudy (anonymous):

can you help me with my unit 6 lesson 11 math if youre that far? conexus 10th grade math

OpenStudy (anonymous):

@DIDITHEDAD

OpenStudy (nikkibxox):

I'm in Connexus 11th. Are you in Geometry or Alg 1?

OpenStudy (anonymous):

Algebra 2 B !+@NikkiBxox

OpenStudy (nikkibxox):

Woahh, Alg 2 already!! Good job. Yeah I'm there. I'll help. What do you need?

OpenStudy (anonymous):

Would you mind if i posted screenshots? @NikkiBxox and im looking for help do to if i don't pass this ill be failing math and get in trouble.:(

OpenStudy (nikkibxox):

Sure thing! The mods might not like it. Is it a test or practice? Or quiz?

OpenStudy (anonymous):

Its a test but i have question 17-24 done i just don't know the answers to 1-16 ive been stressing over it like crazy :( give me one second to get the screen shots @NikkiBxox

OpenStudy (nikkibxox):

Alright. I can help.

OpenStudy (anonymous):

@NikkiBxox these are the 15 questions im stuck on.

OpenStudy (nikkibxox):

Let me tag some people and do the problems I can, Okay? @TheDebunker @RainbowsxD @freckles

OpenStudy (anonymous):

okay @NikkiBxox

OpenStudy (anonymous):

can you help me on mine plz

OpenStudy (anonymous):

@LegendaryClasher @robtobey @Roseamora

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

can u help her trojan

OpenStudy (trojanpoem):

That's combinations and permutations: 1) 7P7 = 5040 ( Arrangement matter) 2) 9!/4! = 1 * 2 * 3 * 4 * 5 * 6 * 7 * 8 * 9 / 4 * 3 * 2 *1 = 15120

OpenStudy (trojanpoem):

The order : 90 caps Green caps:18 Probability it will be green 18 * 100/90 = 20 percent Probability it won't be green: 100 - 20 = 80 percent

OpenStudy (trojanpoem):

I think, I am just answering.

OpenStudy (anonymous):

yah

OpenStudy (trojanpoem):

@BABYKROWW , I don't know how to explain that. Look for someone who is good at explaining.

OpenStudy (anonymous):

ok brb

OpenStudy (trojanpoem):

4) tossed : 50 coins heads :22 TIMES 22/50 = 11/25 5) 28/29

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@Englishguy

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

I have an awesome site for you that will help you on almost any math problem, you just type in the problem and it shows you how to do it want me to send you the link?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm just going to send you the link https://m.mathway.com/

OpenStudy (anonymous):

that doesnt help with these questions ive tried @Englishguy @TrojanPoem You were helping so unless someone else helps ill give you a medal!(:

OpenStudy (anonymous):

Are they story problems?

OpenStudy (anonymous):

ive been tryinggggggg

OpenStudy (anonymous):

@Englishguy most of them are and @BABYKROWW i know i just will probaby open a new question and tag you in a comment just to give you a medal since i cant give two on one post

OpenStudy (anonymous):

im srry i couldn'y help

OpenStudy (anonymous):

I'm going to try these

OpenStudy (anonymous):

I got the first one so far

OpenStudy (anonymous):

i need 6-16 now.

OpenStudy (anonymous):

Number 6 the answer is 37.5%

OpenStudy (anonymous):

@Peach-es1234 can you hep me with 6-16?

OpenStudy (anonymous):

So you round it to 38%

OpenStudy (anonymous):

Can you open a new one So I can see what your question is? :)

OpenStudy (anonymous):

Number 7 the answer is 1/28

OpenStudy (anonymous):

@Englishguy those arent even answer choices and @Peach-es1234 i can (:

OpenStudy (anonymous):

Thanks :)